leila123456:

Michelle and Andrea are walking laps for charity. The amount of money, y, that each girl raises for a certain number of laps, x, is shown in the tables below. Michelle A 2-column table with 3 rows. Column 1 is labeled x with entries 10, 30, 60. Column 2 is labeled y with entries 15, 45, 90. Andrea A 2-column table with 3 rows. Column 1 is labeled x with entries 20, 32, 40. Column 2 is labeled y with entries 25, 40, 50. Which statement explains who will raise $100 with the fewest number of laps? Michelle will, because the slope of the line described by the data in her table is the greatest. Andrea will, because the slope of the line described by the data in her table is the greatest. Michelle will, because her number of laps are increasing at a faster rate. Andrea will, because her number of laps are increasing at a slower rate.

leila123456:

Michelle

1 attachment
leila123456:

Andrea

1 attachment
darkknight:

So first, we need to see the equation for each of the girls

darkknight:

Can you find the slope/ equation for each of them?

darkknight:

Michelle is 10,15 and 30,45 So 45-15)/(30-10) = 30/20 = 1.5 Michelle, y= 1.5x

leila123456:

So do the slope formula?

darkknight:

Essentially yes. Because both equations seem to have a y-intercept of 0. Like you won't make money without running laps

darkknight:

For Andrea 20,25) and (32,40) So 40-25)/(32-20) = 15/12 = 5/4 y=5/4x y=1.25x Now based off of this whichever girl has the higher slope will get 100$ first.

leila123456:

k

darkknight:

So who is it?

leila123456:

Michelle

darkknight:

leila123456:

thx

darkknight:

np